# Time-Scale Coupling Between States and Parameters in Recurrent Neural Networks

### v2

@@ 5.1.2 Single scalar gate
$J^{g}_{j}=\frac{\partial g_{j}}{\partial x_{j}}=\sigma^{\prime}(a^{g}_{j})W^{r,g},\qquad G_{j-1}=d_{j-1}J^{g}_{j-1},\quad d_{j-1}=\phi(a_{j-1})-x_{j-1}.$ (32)
@@
5.1.3 Multiple gates
For the multi-gated model (16), the gate pre-activation is a vector $a^{g}_{j}\in\mathbb{R}^{N_{r}}$, yielding the Jacobian
@@ 5.1.3 Multiple gates
with $D^{g}_{j}=\mathrm{diag}(\sigma^{\prime}(a^{g}_{j}))$. In this case, $g_{j}\in\mathbb{R}^{N_{r}}$ is a vector of unit-wise update rates.
@@
5.2 Time-scale interaction
@@
5.2.1 Constant gate case
In the constant gate case $g_{t}\equiv\alpha\in(0,1]$, the Jacobian (27) is
$J_{j}\;=\;I+\alpha\,A_{j-1},\qquad A_{j-1}:=D_{j-1}W_{r}-I.$ (38)
The BPTT gradient can then be written as
$\frac{\partial E_{t}}{\partial\theta_{l}}=\frac{\partial E_{t}}{\partial x_{t}}\sum_{k=1}^{t}\left[\prod_{j=k+1}^{t}\big{(}I+\alpha\,A_{j-1}\big{)}\right]\frac{\partial x_{k}}{\partial\theta_{l}},$ (39)
where $\theta_{l}$ denotes any parameter of the network (weights, biases, gating parameters, etc.).
Factoring $\alpha$ out of each Jacobian in the product gives
$\prod_{j=k+1}^{t}\big{(}I+\alpha\,A_{j-1}\big{)}\;=\;\left(\prod_{j=k+1}^{t}\alpha\right)\;\prod_{j=k+1}^{t}\big{(}\alpha^{-1}I+A_{j-1}\big{)}\;=\;\alpha^{\,t-k}\;\mathcal{P}_{t,k},$ (40)
where we defined the normalized product
$\mathcal{P}_{t,k}\;:=\;\prod_{j=k+1}^{t}\big{(}\alpha^{-1}I+A_{j-1}\big{)}.$ (41)
@@ 5.2.1 Constant gate case
5.2.2 RNNs with scalar gate
When considering the Jacobian in (31), the product of Jacobians inside the gradient update formula can be rewritten as:
$\prod_{j=k+1}^{t}G_{j-1}+g_{j-1}D_{j-1}W^{r}+(1-g_{j-1})I=\prod_{j=k+1}^{t}G_{j-1}+I+g_{j-1}(D_{j-1}W^{r}-I).$ (44)
@@ 5.2.2 RNNs with scalar gate
$\frac{\partial\mathcal{E}_{t}}{\partial\theta_{l}}=\frac{\partial\mathcal{E}_{t}}{\partial x_{t}}\sum_{k=1}^{t}\left[\prod_{j=k+1}^{t}G_{j-1}+I+g_{j-1}(D_{j-1}W^{r}-I)\right]\frac{\partial x_{k}}{\partial\theta_{l}}.$ (45)
@@
Expanding the matrix product to the first order in the rank-1 corrections $G_{j-1}$, we obtain
$\prod_{j=k+1}^{t}M_{j-1}\approx\underbrace{\prod_{j=k+1}^{t}\big{(}I+g_{j-1}A_{j-1}\big{)}}_{\text{dominant dynamics}}+\sum_{m=k+1}^{t}\Bigg{(}\prod_{j=m+1}^{t}(I+g_{j-1}A_{j-1})\Bigg{)}G_{m-1}\Bigg{(}\prod_{j=k+1}^{m-1}(I+g_{j-1}A_{j-1})\Bigg{)}.$ (47)
The first term describes the main dynamics without the rank-1 updates, while the second term is a sum of rank-1 corrections, each inserted at a different time step $m$ and propagated by the main dynamics. We now factor the gate scalars $g_{j-1}$ explicitly from the first term in (47) like we did in Section 5.2.1.
$\prod_{j=k+1}^{t}(I+g_{j-1}A_{j-1})=\Bigg{(}\prod_{j=k+1}^{t}g_{j-1}\Bigg{)}\Bigg{(}\prod_{j=k+1}^{t}\big{(}g_{j-1}^{-1}I+A_{j-1}\big{)}\Bigg{)}.$
@@
$\prod_{j=k+1}^{t}M_{j-1}\approx\Bigg{(}\prod_{j=k+1}^{t}g_{j-1}\Bigg{)}P_{t,k}+\sum_{m=k+1}^{t}\Bigg{(}\prod_{j=m+1}^{t}g_{j-1}\Bigg{)}R_{t,m}\,G_{m-1}\,\Bigg{(}\prod_{j=k+1}^{m-1}g_{j-1}\Bigg{)}S_{k,m},$ (48)
where
$P_{t,k}=\prod_{j=k+1}^{t}\big{(}g_{j-1}^{-1}I+A_{j-1}\big{)},$
and the normalized matrices $R_{t,m}$ and $S_{k,m}$ are defined as
$R_{t,m}=\prod_{j=m+1}^{t}\big{(}g_{j-1}^{-1}I+A_{j-1}\big{)},\qquad S_{k,m}=\prod_{j=k+1}^{m-1}\big{(}g_{j-1}^{-1}I+A_{j-1}\big{)}.$
Plugging (48) into the gradient expression, we obtain
$\frac{\partial E_{t}}{\partial\theta_{l}}=\frac{\partial E_{t}}{\partial x_{t}}\sum_{k=1}^{t}\left[\left(\prod_{j=k+1}^{t}g_{j-1}\right)P_{t,k}+\sum_{m=k+1}^{t}\left(\prod_{j=m+1}^{t}g_{j-1}\right)R_{t,m}\,G_{m-1}\,\left(\prod_{j=k+1}^{m-1}g_{j-1}\right)S_{k,m}\right]\frac{\partial x_{k}}{\partial\theta_{l}}.$ (49)
@@
Here the perturbative correction terms arise from the residual $(1-g_{t})x_{t}$ pathway and from gate-gradient contributions. These introduce rank-1 modifications to the Jacobian products, modulating gradient propagation in a way that cannot be collapsed into a single scalar factor. Thus, while the multiplicative gate product sets the dominant time-varying decay, the perturbative corrections adjust the effective learning rate in a lag- and state-dependent manner.
5.2.3 RNNs with multiple gates
When considering the Jacobian in (36) the gradient equation becomes:
$\frac{\partial\mathcal{E}_{t}}{\partial\theta_{l}}=\frac{\partial\mathcal{E}_{t}}{\partial x_{t}}\sum_{k=1}^{t}\left[\prod_{j=k+1}^{t}G_{j-1}+\left(\hat{g}_{j-1}D_{j-1}W^{r}\right)+\left(\hat{\bar{g}}_{j-1}I\right)\right]\frac{\partial x_{k}}{\partial\theta_{l}}.$ (51)
@@ 5.2.3 RNNs with multiple gates
where
$\tilde{A}_{j-1}=A_{j-1}+\Gamma_{j-1}^{-1}(I-\Gamma_{j-1}),$
with $\Gamma_{j-1}^{-1}(I-\Gamma_{j-1})$ defined element-wise as $\mathrm{diag}((1-g^{(i)}_{j-1})/g^{(i)}_{j-1})$ for $g^{(i)}_{j-1}>0$.
@@
$\prod_{j=k+1}^{t}M_{j-1}\approx\underbrace{\prod_{j=k+1}^{t}\big{(}\Gamma_{j-1}\tilde{A}_{j-1}\big{)}}_{\text{dominant dynamics}}+\sum_{m=k+1}^{t}\Bigg{(}\prod_{j=m+1}^{t}\Gamma_{j-1}\tilde{A}_{j-1}\Bigg{)}G_{m-1}\Bigg{(}\prod_{j=k+1}^{m-1}\Gamma_{j-1}\tilde{A}_{j-1}\Bigg{)}.$ (54)
@@
where
$P_{t,k}=\prod_{j=k+1}^{t}\tilde{A}_{j-1},$
and the product $\prod_{j=k+1}^{t}\Gamma_{j-1}$ is a diagonal matrix with entries
$\left(\prod_{j=k+1}^{t}g^{(1)}_{j-1},\ldots,\prod_{j=k+1}^{t}g^{(N_{r})}_{j-1}\right)$.
@@
The full expansion to the first order becomes
$\prod_{j=k+1}^{t}M_{j-1}\approx\Bigg{(}\prod_{j=k+1}^{t}\Gamma_{j-1}\Bigg{)}P_{t,k}+\sum_{m=k+1}^{t}\Bigg{(}\prod_{j=m+1}^{t}\Gamma_{j-1}\Bigg{)}R_{t,m}\,G_{m-1}\,\Bigg{(}\prod_{j=k+1}^{m-1}\Gamma_{j-1}\Bigg{)}S_{k,m}.$ (56)
@@
$\frac{\partial E_{t}}{\partial\theta_{l}}=\frac{\partial E_{t}}{\partial x_{t}}\sum_{k=1}^{t}\left[\left(\prod_{j=k+1}^{t}\Gamma_{j-1}\right)P_{t,k}+\sum_{m=k+1}^{t}\left(\prod_{j=m+1}^{t}\Gamma_{j-1}\right)R_{t,m}\,G_{m-1}\,\left(\prod_{j=k+1}^{m-1}\Gamma_{j-1}\right)S_{k,m}\right]\frac{\partial x_{k}}{\partial\theta_{l}}.$ (57)
@@
Here the perturbative correction terms involve interactions between gates and state updates, producing full-rank contributions that cannot be simplified into a scalar or diagonal factor. Their effect is modulated by the diagonal gate products but introduces additional couplings across neurons. In the multi-gate case:
  • Each neuron has its own time-varying effective learning rate determined by its specific gate trajectory.
  • Gradient propagation is anisotropic: some directions are damped strongly, while others are preserved, depending on the distribution of gate trajectories across neurons.
  • The correction terms act as neuron-coupling perturbations, making the effective scaling both magnitude- and direction-dependent.
6 Analogies between gating mechanisms and adaptive gradient descent methods
We now turn to well-known variations of the basic gradient descent algorithm [27]. Our goal is to draw parallels between these algorithmic modifications and the implicit effects induced by tunable time scales in RNN state dynamics. In particular, we examine how gates – by shaping the Jacobian structure – alter the effective optimization behaviour of plain gradient descent, in ways reminiscent of learning rate schedules, momentum terms, or adaptive update rules.
The analysis in Sections 5.2.15.2.3 shows that the Jacobian factors in gated RNNs naturally lead to multiplicative modulation of the backpropagated gradient. From the first-order expansion, this modulation directly translates into an effective learning rate that depends on the gate configuration. This behavior closely parallels well-known adaptive variants of gradient descent.
Constant gate ($\alpha$).
@@ 6 Analogies between gating mechanisms and adaptive gradient descent methods
Single time-varying gate ($g_{j-1}$).
@@
Table 1: Conceptual similarities between gating-induced gradient modulation and adaptive gradient descent methods.
In summary, the presence of gates in RNN dynamics not only stabilizes gradient propagation across long sequences but also induces implicit, data-driven modulation of the effective learning rate. Depending on the gating configuration, this modulation can behave like a fixed preconditioner, a dynamic learning rate schedule, or a fully adaptive per-parameter step size—mirroring the design principles of modern optimization algorithms without explicitly modifying the optimizer.
7 Simulations
7.1 Effective learning rate induced by gates
We empirically verify that, even when SGD uses a fixed global step size $\mu$, gating mechanisms induce a lag-dependent effective learning rate that depends on the state-space model (leaky, scalar gate, multi-gate). The central object we use is the lag-conditioned sensitivity
$S_{t,k}(\theta)\;=\;\Big{\|}\,\prod_{j=k+1}^{t}J_{j}(\theta)\,\Big{\|}_{2},$
@@ 7.1 Effective learning rate induced by gates
$\tilde{\mu}_{\mathrm{eff}}(h;\ell)\;\propto\;\mathrm{median}_{(t,k):\,t-k=h}\,S_{t,k}(\theta_{\ell}).$
@@
$\tilde{\mu}_{\mathrm{eff}}(h;\ell)\;=\;\frac{\mathrm{median}_{(t,k):\,h}\,S_{t,k}(\theta_{\ell})}{\mathrm{median}_{(t,k):\,h=1}\,S_{t,k}(\theta_{\ell})}\,.$
@@
Predictors and overlays.
To interpret the empirical effective learning rates, we compare them against a simple gate-product predictor derived from the zeroth-order theory of Section 5.2. For each model we define
$P_{t,k}(\theta_{\ell})\;=\;\begin{cases}\alpha^{\,t-k},&\text{leaky (constant $\alpha$)},\\[2.0pt] \prod_{j=k+1}^{t}g_{j-1},&\text{scalar gate},\\[6.0pt] \frac{1}{N_{r}}\displaystyle\sum_{i=1}^{N_{r}}\;\prod_{j=k+1}^{t}g^{(i)}_{j-1},&\text{multi-gate},\end{cases}$
@@
$\bar{P}(h;\ell)=\mathrm{median}_{(t,k):\,t-k=h}\,P_{t,k}(\theta_{\ell}).$
@@
$\tilde{\mu}^{(0)}_{\mathrm{pred}}(h;\ell)$ $=\frac{\bar{P}(h;\ell)}{\bar{P}(1;\ell)},$ (zeroth-order theory, no fitting)
$\tilde{\mu}^{(\mathrm{fit})}_{\mathrm{pred}}(h;\ell)$ $=\Bigg{(}\frac{\bar{P}(h;\ell)}{\bar{P}(1;\ell)}\Bigg{)}^{s(\ell)},$ $\text{(empirical correction)}.$
@@
Setup.
We train each model (leaky, scalar, multi-gate) with plain SGD (no momentum, no adaptivity) on a simple sequence task – adding problem. At selected checkpoints, we (i) run a fixed probe batch, (ii) build the exact one-step Jacobians $J_{j}$ from Section 5.1, (iii) estimate $S_{t,k}$ via a short power method iteration (5 steps) for the products $\prod_{j=k+1}^{t}J_{j}$, a standard technique to approximate the largest singular value, (iv) compute $P_{t,k}$, and (v) produce the normalized lag profiles with theory overlays. We also track the fitted slope $s(\ell)$ across training.
Results.
Across all models, the empirical profiles $\tilde{\mu}_{\mathrm{eff}}(h;\ell)$ exhibit a clear, monotone attenuation with lag and are well tracked by the fitted-power overlay $\tilde{\mu}^{(\mathrm{fit})}_{\mathrm{pred}}(h;\ell)$; the zeroth-order overlay $\tilde{\mu}^{(0)}_{\mathrm{pred}}(h;\ell)$ captures the right trend but is often too conservative or too aggressive depending on the model. The fitted log–log slopes extracted from the summaries (slope, $R^{2}$) are:

### v3

@@ V-A2 Single scalar gate
$J^{g}_{j}=\frac{\partial g_{j}}{\partial x_{j}}=\sigma^{\prime}(a^{g}_{j})W^{r,g},\qquad G_{j-1}=d_{j-1}J^{g}_{j-1},$ (19)
$\quad d_{j-1}=\phi(a_{j-1})-x_{j-1}.$
@@
V-A3 Multiple gates
For the multi-gated model (12), the gate pre-activation is a vector $a^{g}_{j}\in\mathbb{R}^{N_{r}}$, yielding the gate Jacobian
@@ V-A3 Multiple gates
with $D^{g}_{j}=\mathrm{diag}(\sigma^{\prime}(a^{g}_{j}))$. In this case, $g_{j}\in\mathbb{R}^{N_{r}}$ is a vector of neuron-wise update rates.
@@
V-B Time-scale coupling
@@
V-B1 Constant gate case
In the constant gate case $g_{t}\equiv\alpha\in(0,1]$, the BPTT gradient can be written as
$\frac{\partial E_{t}}{\partial\theta_{l}}=\frac{\partial E_{t}}{\partial x_{t}}\sum_{k=1}^{t}\left[\prod_{j=k+1}^{t}\big{(}I+\alpha\,A_{j-1}\big{)}\right]\frac{\partial x_{k}}{\partial\theta_{l}}.$ (23)
where the Jacobian (17) is $J_{j}\;=\;I+\alpha\,A_{j-1},\ A_{j-1}:=D_{j-1}W_{r}-I$. Factoring $\alpha$ out of each Jacobian in the product gives
$\prod_{j=k+1}^{t}\big{(}I+\alpha\,A_{j-1}\big{)}$ (24)
$\;=\;\left(\prod_{j=k+1}^{t}\alpha\right)\;\prod_{j=k+1}^{t}\big{(}\alpha^{-1}I+A_{j-1}\big{)}\;=\;\alpha^{\,t-k}\;\mathcal{P}_{t,k},$ (25)
where we defined the normalized product $\mathcal{P}_{t,k}\;:=\;\prod_{j=k+1}^{t}\big{(}\alpha^{-1}I+A_{j-1}\big{)}$.
@@ V-B1 Constant gate case
V-B2 RNNs with scalar gate
When considering the Jacobian in (LABEL:eq:jacobian_rnn_dt_warping), the product of Jacobians inside the gradient update formula can be rewritten as:
$\prod_{j=k+1}^{t}G_{j-1}+g_{j-1}D_{j-1}W^{r}+(1-g_{j-1})I$ (28)
$=\prod_{j=k+1}^{t}G_{j-1}+I+g_{j-1}(D_{j-1}W^{r}-I).$
@@ V-B2 RNNs with scalar gate
$\frac{\partial\mathcal{E}_{t}}{\partial\theta_{l}}$ (29)
$=\frac{\partial\mathcal{E}_{t}}{\partial x_{t}}\sum_{k=1}^{t}\left[\prod_{j=k+1}^{t}G_{j-1}+I+g_{j-1}(D_{j-1}W^{r}-I)\right]\frac{\partial x_{k}}{\partial\theta_{l}}.$
@@
Expanding the matrix product (LABEL:eq:jacobian_product_scalar_gate) to the first order in the rank-1 corrections $G_{j-1}$, we obtain
$\prod_{j=k+1}^{t}M_{j-1}$ $\approx\prod_{j=k+1}^{t}\big{(}I+g_{j-1}A_{j-1}\big{)}$ (31)
$+\sum_{m=k+1}^{t}$ $\Bigg{(}\prod_{j=m+1}^{t}(I+g_{j-1}A_{j-1})\Bigg{)}$
$G_{m-1}\Bigg{(}\prod_{j=k+1}^{m-1}(I+g_{j-1}A_{j-1})\Bigg{)}.$
The first term describes the main dynamics without the rank-1 updates, while the second term is a sum of rank-1 perturbative corrections, each inserted at a different time step $m$ and propagated by the main dynamics. We now factor the gate scalars $g_{j-1}$ explicitly from the first term in (31) like we did in Section V-B1.
$\prod_{j=k+1}^{t}(I+g_{j-1}A_{j-1})=$ (32)
$\Bigg{(}\prod_{j=k+1}^{t}g_{j-1}\Bigg{)}\Bigg{(}\prod_{j=k+1}^{t}\big{(}g_{j-1}^{-1}I+A_{j-1}\big{)}\Bigg{)}.$
@@
$\prod_{j=k+1}^{t}M_{j-1}\approx\Bigg{(}\prod_{j=k+1}^{t}g_{j-1}\Bigg{)}P_{t,k}$ (33)
$+\sum_{m=k+1}^{t}\Bigg{(}\prod_{j=m+1}^{t}g_{j-1}\Bigg{)}R_{t,m}G_{m-1}\,\Bigg{(}\prod_{j=k+1}^{m-1}g_{j-1}\Bigg{)}S_{k,m},$
where $P_{t,k}=\prod_{j=k+1}^{t}\big{(}g_{j-1}^{-1}I+A_{j-1}\big{)}$ and the normalized matrices are defined as $R_{t,m}=\prod_{j=m+1}^{t}\big{(}g_{j-1}^{-1}I+A_{j-1}\big{)}$ and $S_{k,m}=\prod_{j=k+1}^{m-1}\big{(}g_{j-1}^{-1}I+A_{j-1}\big{)}$.
Plugging (LABEL:eq:single_gate-product-decomposition) into the gradient expression, we obtain
$\frac{\partial E_{t}}{\partial\theta_{l}}$ $=\frac{\partial E_{t}}{\partial x_{t}}\sum_{k=1}^{t}\left(\prod_{j=k+1}^{t}g_{j-1}\right)P_{t,k}$ (34)
$+\sum_{m=k+1}^{t}\left(\prod_{j=m+1}^{t}g_{j-1}\right)R_{t,m}$
$G_{m-1}\,\left(\prod_{j=k+1}^{m-1}g_{j-1}\right)S_{k,m}\frac{\partial x_{k}}{\partial\theta_{l}}.$
@@
Here the perturbative correction terms arise from the residual $(1-g_{t})x_{t}$ pathway and from gate-gradient contributions. These introduce rank-1 modifications to the Jacobian products, modulating gradient propagation in a way that cannot be collapsed into a single scalar factor.
V-B3 RNNs with multiple gates
When considering the Jacobian in (LABEL:eq:jacobian_rnn_dt_multiple_warping) the gradient equation becomes:
$\frac{\partial\mathcal{E}_{t}}{\partial\theta_{l}}$ (36)
$=\frac{\partial\mathcal{E}_{t}}{\partial x_{t}}\sum_{k=1}^{t}\left[\prod_{j=k+1}^{t}G_{j-1}+\left(\hat{g}_{j-1}D_{j-1}W^{r}\right)+\left(\hat{\bar{g}}_{j-1}I\right)\right]\frac{\partial x_{k}}{\partial\theta_{l}}.$
@@ V-B3 RNNs with multiple gates
where $\tilde{A}_{j-1}=A_{j-1}+\Gamma_{j-1}^{-1}(I-\Gamma_{j-1})$, with $\Gamma_{j-1}^{-1}(I-\Gamma_{j-1})$ defined element-wise as $\mathrm{diag}((1-g^{(i)}_{j-1})/g^{(i)}_{j-1})$ for $g^{(i)}_{j-1}>0$.
@@
$\prod_{j=k+1}^{t}M_{j-1}\approx\prod_{j=k+1}^{t}\big{(}\Gamma_{j-1}\tilde{A}_{j-1}\big{)}$ (39)
$+\sum_{m=k+1}^{t}\Bigg{(}\prod_{j=m+1}^{t}\Gamma_{j-1}\tilde{A}_{j-1}\Bigg{)}G_{m-1}\Bigg{(}\prod_{j=k+1}^{m-1}\Gamma_{j-1}\tilde{A}_{j-1}\Bigg{)}.$
@@
where $P_{t,k}=\prod_{j=k+1}^{t}\tilde{A}_{j-1}$, and the product $\prod_{j=k+1}^{t}\Gamma_{j-1}$ is a diagonal matrix with entries $\left(\prod_{j=k+1}^{t}g^{(1)}_{j-1},\ldots,\prod_{j=k+1}^{t}g^{(N_{r})}_{j-1}\right)$.
@@
The expansion to the first order becomes
$\prod_{j=k+1}^{t}M_{j-1}\approx\Bigg{(}\prod_{j=k+1}^{t}\Gamma_{j-1}\Bigg{)}P_{t,k}$ (41)
$+\sum_{m=k+1}^{t}\Bigg{(}\prod_{j=m+1}^{t}\Gamma_{j-1}\Bigg{)}R_{t,m}G_{m-1}\,\Bigg{(}\prod_{j=k+1}^{m-1}\Gamma_{j-1}\Bigg{)}S_{k,m}.$
@@
$\frac{\partial E_{t}}{\partial\theta_{l}}$ $=\frac{\partial E_{t}}{\partial x_{t}}\sum_{k=1}^{t}\left(\prod_{j=k+1}^{t}\Gamma_{j-1}\right)P_{t,k}$ (42)
$+\sum_{m=k+1}^{t}\left(\prod_{j=m+1}^{t}\Gamma_{j-1}\right)R_{t,m}$
$G_{m-1}\,\left(\prod_{j=k+1}^{m-1}\Gamma_{j-1}\right)S_{k,m}\frac{\partial x_{k}}{\partial\theta_{l}}.$
@@
Here the perturbative correction terms involve interactions between gates and state updates, producing full-rank contributions that cannot be simplified into a scalar or diagonal factor. In the multi-gate case: (i) each neuron has its own time-varying effective learning rate determined by its specific gate trajectory; (ii) gradient propagation is substantially anisotropic: some directions may be damped strongly, while others may be preserved, depending on the distribution of gate trajectories across neurons; (iii) the correction terms act as neuron-coupling perturbations, making the effective scaling both magnitude- and direction-dependent.
VI Analogies between gating mechanisms and adaptive gradient descent methods
We now turn to well-known variations of the basic gradient descent algorithm [32]. Our goal is to draw parallels between these algorithmic modifications and the implicit effects induced by tunable time scales in RNN state dynamics. A summary is shown in Table I. In particular, we examine how gates – by shaping the Jacobian structure – alter the effective optimization behaviour of plain gradient descent, in ways reminiscent of learning rate schedules, momentum terms, or adaptive update rules.
The analysis in Sections V-B1V-B3 shows that the Jacobian factors in gated RNNs naturally lead to multiplicative modulation of the backpropagated gradient. From the first-order expansion, this modulation directly translates into an effective learning rate that depends on the gate configuration.
Constant gate ($\alpha$)
@@ VI Analogies between gating mechanisms and adaptive gradient descent methods
Single time-varying gate ($g_{j-1}$)
@@
TABLE I: Conceptual similarities between gating-induced gradient modulation and adaptive gradient descent methods.
@@
VII Simulations
VII-A Effective learning rate induced by gates
We empirically verify that, even when SGD uses a fixed global step size $\mu$, gating mechanisms induce a lag-dependent effective learning rate that depends on the state-space model. The central object we use is the lag-conditioned sensitivity
$S_{t,k}\;=\;\left\|\,\prod_{j=k+1}^{t}J_{j}\,\right\|_{2},$ (44)
@@ VII-A Effective learning rate induced by gates
$\tilde{\mu}_{\mathrm{eff}}(h;\ell)\;\propto\;\mathrm{median}_{h}\,S_{t,k}(\theta_{\ell}).$ (45)
@@
$\tilde{\mu}_{\mathrm{eff}}(h;\ell)\;=\;\frac{\mathrm{median}_{h}\,S_{t,k}(\theta_{\ell})}{\mathrm{median}_{h=1}\,S_{t,k}(\theta_{\ell})}\,.$ (46)
@@
Predictors and overlays
To interpret the empirical effective learning rates, we compare them against a simple gate-product predictor derived from the zeroth-order theory of Section V-B. For each model we define
$P_{t,k}(\theta_{\ell})\;=\;\begin{cases}\alpha^{\,t-k},&\text{leaky (constant $\alpha$)},\\[2.0pt] \prod_{j=k+1}^{t}g_{j-1},&\text{scalar gate},\\[6.0pt] \frac{1}{N_{r}}\displaystyle\sum_{i=1}^{N_{r}}\;\prod_{j=k+1}^{t}g^{(i)}_{j-1},&\text{multi-gate},\end{cases}$ (47)
@@
$\bar{P}(h;\ell)=\mathrm{median}_{h}\,P_{t,k}(\theta_{\ell}).$ (48)
@@
$\tilde{\mu}^{(0)}_{\mathrm{pred}}(h;\ell)$ $=\frac{\bar{P}(h;\ell)}{\bar{P}(1;\ell)},$ (zeroth-order theory, no fitting) (49)
$\tilde{\mu}^{(\mathrm{fit})}_{\mathrm{pred}}(h;\ell)$ $=\Bigg{(}\frac{\bar{P}(h;\ell)}{\bar{P}(1;\ell)}\Bigg{)}^{s(\ell)},$ $\text{(empirical correction)}.$ (50)
@@
Interpretation of the slope
The slope provides a quantitative diagnostic of how gates renormalize effective learning rates:
  • $s(\ell)=0$: no lag dependence, i.e. the effective learning rate remains constant and equal to the nominal $\mu$.
  • $s(\ell)=1$: exact agreement with the zeroth-order gate-product predictor, i.e. $\tilde{\mu}_{\mathrm{eff}}(h)\propto P_{t,k}$.
  • $s(\ell)>1$: effective sensitivity decays more steeply with lag than the predictor, shortening the memory horizon and yielding more conservative updates.
  • $0<s(\ell)<1$: effective sensitivity decays more slowly with lag than the predictor, extending the memory horizon and yielding more permissive updates.
Setup
We train each RNN model (leaky (8), scalar (11), multi-gate (12)) with plain SGD (no momentum, no adaptivity) on a simple sequence task – adding problem. At selected checkpoints, we (i) run a fixed probe batch, (ii) build the exact one-step Jacobians $J_{j}$ from Section V-A, (iii) estimate $S_{t,k}$ via a short power method iteration (5 steps) for the products $\prod_{j=k+1}^{t}J_{j}$, a standard technique to approximate the largest singular value, (iv) compute $P_{t,k}$, and (v) produce the normalized lag profiles with theory overlays. We also track the fitted slope $s(\ell)$ across training.
Results
Results are shown in Figures 911. Across all models, the empirical profiles $\tilde{\mu}_{\mathrm{eff}}(h;\ell)$ exhibit a clear, monotone attenuation with lag and are well tracked by the fitted-power overlay $\tilde{\mu}^{(\mathrm{fit})}_{\mathrm{pred}}(h;\ell)$; the zeroth-order overlay $\tilde{\mu}^{(0)}_{\mathrm{pred}}(h;\ell)$ captures the right trend but is often too conservative or too aggressive depending on the model. The fitted log–log slopes extracted from the summaries (slope, $R^{2}$) are: